# KG-IRAG: A Knowledge Graph-Based Iterative Retrieval-Augmented Generation Framework for Temporal Reasoning

### v2

@@ Abstract
KG-IRAG: A Knowledge Graph-Based Iterative Retrieval-Augmented Generation Framework for Temporal Reasoning
Ruiyi Yang1  Hao Xue1  Imran Razzak2,1  Hakim Hacid3  Flora D. Salim1 1University of New South Wales, Sydney, NSW, Austalia 2Mohamed Bin Zayed University of Artificial Intelligence, Abu Dhabi, UAE 3Technology Innovation Institute, Abu Dhabi, UAE {ruiyi.yang, hao.xue1, flora.salim}@unsw.edu.au imran.razzak@mbzuai.ac.ae, hakim.hacid@tii.ae
@@ 1 Introduction
Knowledge Graphs (KGs) represent entities, their attributes, and relationships in a structured form, often employed to facilitate information retrieval, recommendation systems, and question answering. By encoding real-world facts as triples (e.g., (Sydney Opera House-[located in]-Sydney)), KGs capture intricate relationships between entities, allowing for more contextually rich predictions. These capabilities have been widely utilized in tasks such as link prediction and knowledge graph completion (KGC). Conventional methods, including TransE Bordes et al. (2013), compute embeddings for entities and relationships to enhance the comprehensiveness of KGs, supporting their use in diverse applications, such as directly information retrieval Dietz et al. (2018), or logical problems like question answering Huang et al. (2019).
@@
Recent advances in Retrieval-Augmented Generation (RAG) have highlighted the potential of integrating external knowledge sources like KGs with large language models (LLMs) to enhance their performance in answering queries that require domain-specific information Zhang et al. (2024a); Siriwardhana et al. (2023). GraphRAG, a variant of RAG, utilizes KGs to enhance the retrieval process through relational data, thus offering more precise and contextually relevant responses from LLMs Wu et al. (2024); Matsumoto et al. (2024); Edge et al. (2024). However, despite the effectiveness of RAG methods in improving retrieval, they often fail to address the challenges posed in time-related tasks, such as planning trips or making decisions based on weather or traffic conditions. In those scenarios, LLMs must retrieve relevant information and reason, involving temporal relationships and require dynamic, context-aware reasoning that goes beyond static fact retrieval. LLMs are limited in their capacity to do such complex reasoning on huge amount of data input without substantial model fine-tuning, and existing RAG methods primarily focus on one-time knowledge retrieval rather than iterative retrieval. Figure 1 shows a scenario in which LLMs tend to generate hallucinations facing trip planning.
To address these limitations, this paper introduces a novel framework, Knowledge Graph-Based Iterative Retrieval-Augmented Generation (KG-IRAG). KG-IRAG integrates KGs with an iterative reasoning process, enabling LLMs to solve problems by incrementally retrieving relevant data through iterative queries and achieve step-by-step reasoning, allowing LLMs to address queries involving temporal dependencies without model fine-tuning. This approach is particularly suited for scenarios where reasoning must be applied alongside dynamic data retrieval. In addition to presenting the KG-IRAG framework, this paper introduces three new datasets: weatherQA-Irish, weatherQA-Sydney and trafficQA-TFNSW. Those datasets are designed to test LLM’s ability to answer queries that require both retrieving uncertain length of temporal information and doing mathematical reasoning, such as determining the best time to reschedule a trip. The proposed KG-IRAG framework is validated through experiments on temporal-related question answering tasks, demonstrating its ability to significantly enhance LLM performance. By leveraging the rich relational structure of KGs and incorporating iterative reasoning, KG-IRAG offers a robust and efficient solution for addressing complex spatial temporal queries.
@@ 2.1 Combination of Graphs with LLMs
Recent research has increasingly focused on combining LLMs with graphs to achieve mutual enhancement. The combination of LLMs with Graph Neural Networks (GNNs) has been shown to significantly improve the modeling capabilities of graph-structured data. LLMs have been shown to contribute to knowledge graph completion, aiding in downstream tasks such as node classification Chen et al. (2023) and link prediction Shu et al. (2024). Additionally, LLMs play a pivotal role in knowledge graph creation, transforming source texts into graphs Edge et al. (2024); Zhang et al. (2024b).
LLMs also enhance performance in Knowledge Graph-based Question Answering (KBQA) tasks, which leverage external knowledge bases to answer user queries Cui et al. (2019); Wu et al. (2019); Fu et al. (2020). KBQA has been applied across various domains, including text understanding and fact-checking Chen et al. (2019); Suresh et al. (2024). Approaches to KBQA are generally categorized into two types: Information Retrieval (IR)-based and Semantic Parsing (SP)-based. IR-based methods directly retrieve information from KG databases and use the returned knowledge to generate answers Jiang et al. (2022, 2024), whereas SP-based methods generate logical forms for queries, which are then used for knowledge retrieval Chakraborty (2024); Fang et al. (2024). Advanced techniques such as Chain of Knowledge (CoK) Li et al. (2023), G-Retriever He et al. (2024), and Chain of Explorations Sanmartin (2024) have been developed to enhance the precision and efficiency of data retrieval from KGs.
@@ 2.2 Retrieval-Augmented Generation (RAG)
RAG enhances the capabilities of LLMs by integrating external knowledge sources during the response generation process. Unlike traditional LLMs, which rely solely on their pre-trained knowledge, RAG enables models to access real-time or domain-specific information from external databases and knowledge sources. Recent studies have explored RAG from various perspectives, including the modalities of databases Zhao et al. (2024), model architectures, training strategies Fan et al. (2024), and the diverse applications of RAG Gao et al. (2023). Effective evaluation of RAG systems requires attention to both the accuracy of knowledge retrieval and the quality of the generated responses Yu et al. (2024).
@@ 2.3 LLMs Temporal Reasoning
Recent advancements have increasingly focused on improving the temporal reasoning capabilities of LLMs. Temporal reasoning in natural language processing (NLP) typically falls into three categories: temporal expression detection and normalization, temporal relation extraction, and event forecasting Yuan et al. (2024). The integration of temporal graphs has enabled LLMs to perform more effectively in tasks such as time comparison Xiong et al. (2024). Several temporal QA datasets have been developed to test LLMs’ temporal reasoning abilities, including TEMPLAMA Dhingra et al. (2022); Tan et al. (2023), TemporalWiki Jang et al. (2022), and time-sensitive QA datasets Chen et al. (2021). By combining LLMs with temporal KGs Lee et al. (2023); Yuan et al. (2024); Xia et al. (2024), more accurate event forecasting has become possible.
@@
3 Methodology
This section describes the iterative process employed by the KG-IRAG framework, detailing the key components that enable the extraction of temporal and spatial data from a knowledge graph (KG) and the logical reasoning required to generate answers. The framework is designed to handle questions that require multi-step reasoning based on temporal and logical comparison.
Algorithm 1 Start process from LLM1
1:Question, LLM1 parameter setting
2:ReasoningPrompt $RP$, Starting time and location $time_{0},location_{0}$
3:$RP\leftarrow\text{LLM1Reasoning}(Question)$
@@ 3 Methodology
5:Starting plan $time_{0},location_{0}$
6:Iterative Retrieval for LLM2($time_{0},location_{0},RP$)
3.1 Preliminaries and method overview
Two LLMs, LLM1 and LLM2, collaborate throughout this process. LLM1 is responsible for two main tasks: (1) identifying the initial plan for KG exploration, specifically the initial start time 0 and location 0, and (2) generating a reasoning prompt that specifies the information required to answer the query and explains its relevance. The output of LLM1, as outlined in Algorithm 1, is passed to LLM2. LLM2 is responsible for evaluating whether the retrieved data, combined with the reasoning prompt, is sufficient to resolve the query or if further retrieval steps are required.
Figure 2 illustrates the iterative RAG process. The methodology proceeds as follows: after identifying the starting time and location, KG exploration is performed iteratively to retrieve relevant triplets. Each iteration yields new data, which is evaluated by LLM2 to determine whether the problem can be solved with the current triplets and reasoning. If the answer remains unresolved, the framework adjusts the search criteria—moving to a different time or location—and continues retrieving new triplets until the answer is generated.
Figure 3: Knowledge graph schema for KG-IRAG
3.2 Knowledge Graph Construction
When transforming raw data into a KG, time, location, and event status (such as rainfall or traffic volume) serve as key entities. Relationships between these entities capture temporal, spatial, and event-based correlations, allowing for structured exploration during query resolution. The construction of the KG ensures that both spatial and temporal dimensions are embedded within the graph, with time being treated as an entity for easier retrieval and reasoning. Each KG provides a foundation for multi-step reasoning, as queries often span across multiple time points and locations. Specifically, locations and time points are modeled as entities, while attributes such as rain status and traffic volume are presented as relations connecting entities. An example KG schema is shown in figure 3.
Algorithm 2 Iterative Retrieval for LLM2
1:$time_{i}$, $loc_{i}$, ReasoningPrompt $RP$, LLM2 parameter setting
2:Final answer after KG-IRAG process
3:Declare
4:     $k=\text{knowledge}$
5:     $\text{CombinePrompt}(\cdot)=\text{CP}$

### v3

# Beyond Single Pass, Looping Through Time: KG-IRAG with Iterative Knowledge Retrieval

@@ Abstract
Beyond Single Pass, Looping Through Time: KG-IRAG with Iterative Knowledge Retrieval
Ruiyi Yang1  Hao Xue1  Imran Razzak2,1  Hakim Hacid3  Flora D. Salim1 1University of New South Wales, Sydney, NSW, Austalia 2Mohamed Bin Zayed University of Artificial Intelligence, Abu Dhabi, UAE 3Technology Innovation Institute, Abu Dhabi, UAE ruiyi.yang@student.unsw.edu.au, hao.xue1@unsw.edu.au imran.razzak@mbzuai.ac.ae, hakim.hacid@tii.ae flora.salim@unsw.edu.au
@@ 1 Introduction
While State-of-the-art LLMs still suffered from limitations like fabricating fact or giving outdated answers for questions without answer or require latest data Kandpal et al. (2023); Zhang et al. (2023), recent advances in Retrieval-Augmented Generation (RAG) offered a solution by integrating external knowledge sources like KGs with large language models (LLMs), the reliable and flexible supplementary knowledge helps LLMs generate more accurate replies to those queries that require domain-specific information Zhang et al. (2024a); Siriwardhana et al. (2023), without extra time for fine-tuning. As a variant of RAG, GraphRAG utilizes KGs to enhance the retrieval process through relational data. KGs capture intricate relationships between entities, allowing for more contextually rich predictions, thus offering more precise and contextually relevant responses from LLMs Wu et al. (2024); Matsumoto et al. (2024); Edge et al. (2024). Although RAG achieves effectiveness in improving retrieval, they often fail to address the challenges posed in time-related tasks, such as planning trips or making decisions based on weather or traffic conditions. Dealing those queries, current RAG framework still faces several challenges:
$\mathbf{C_{1}}:\textbf{Accurate Retrieval}$ LLMs need to generate a clear data retrieving plan, avoiding both insufficient(wrong answer) and redundant(waste of token) data recall Cuconasu et al. (2024).
$\mathbf{C_{2}}:\textbf{Hallucination}$ Even with correct extra data input, facing too many data in prompts, especially numbers, LLMs tend to generate hallucinations and not following correct facts Gao et al. (2023a).
[…]
Figure 1 shows a scenario in which LLMs tend to generate hallucinations facing trip planning. To do such complex reasoning on huge amount of data input, existing RAG methods primarily focus on one-time knowledge retrieval rather than iterative retrieval. To address these limitations, this paper introduces a novel framework, Knowledge Graph-Based Iterative Retrieval-Augmented Generation (KG-IRAG). KG-IRAG integrates KGs with an iterative reasoning process, enabling LLMs to solve problems by incrementally retrieving relevant data through multi-step queries and achieve step-by-step reasoning. The proposed KG-IRAG framework is validated through experiments on datasets containing queries involving temporal dependencies, demonstrating its ability to significantly enhance LLM performance. Specifically, our key contributions are:
  • Iterative RAG Framework: The new framework enables LLM reasoning in time-sensitive KG subgraphs through cyclic query refinement, offering a robust and efficient solution for addressing complex spatial temporal queries. The context-aware information retrieval via graph walks mirrors human-like reasoning patterns.
  • New datasets targeting complex temporal queries: weatherQA-Irish, weatherQA-Sydney and trafficQA-TFNSW are designed to test LLM’s ability to answer queries that require both retrieving uncertain length of temporal information and doing mathematical reasoning.
@@ 2.1 Combination of Graphs with LLMs
The combinations of LLMs and graphs achieves mutual enhancement. The combination of LLMs with Graph Neural Networks (GNNs) has been shown to significantly improve the modeling capabilities of graph-structured data. LLMs have been shown to contribute to knowledge graph completion, aiding in downstream tasks such as node classification Chen et al. (2023) and link prediction Shu et al. (2024). Additionally, LLMs play a pivotal role in knowledge graph creation, transforming source texts into graphs Edge et al. (2024); Zhang et al. (2024b). LLMs also enhance performance in Knowledge Graph-based Question Answering (KBQA) tasks, which leverage external knowledge bases to answer user queries Cui et al. (2019); Wu et al. (2019); Fu et al. (2020). KBQA tasks includes text understanding and fact-checking Chen et al. (2019); Suresh et al. (2024), and approaches are generally categorized into two types: Information Retrieval (IR)-based and Semantic Parsing (SP)-based. IR-based methods directly retrieve information from KG databases and use the returned knowledge to generate answers Jiang et al. (2022, 2024), whereas SP-based methods generate logical forms for queries, which are then used for knowledge retrieval Chakraborty (2024); Fang et al. (2024). Advanced techniques such as Chain of Knowledge (CoK) Li et al. (2023), G-Retriever He et al. (2024), and Chain of Explorations Sanmartin (2024) have been developed to enhance the precision and efficiency of data retrieval from KGs.
@@ 2.2 Retrieval-Augmented Generation (RAG)
RAG enhances the capabilities of LLMs by integrating external knowledge sources during the response generation process. Unlike traditional LLMs, which rely solely on their pre-trained knowledge, RAG enables models to access real-time or domain-specific information from external databases and knowledge sources. Recent studies have explored RAG from various perspectives, including the modalities of databases Zhao et al. (2024), model architectures, training strategies Fan et al. (2024), and the diverse applications of RAG Gao et al. (2023b). Effective evaluation of RAG systems requires attention to both the accuracy of knowledge retrieval and the quality of the generated responses Yu et al. (2024).
@@ 2.3 LLMs Temporal Reasoning
Recent advancements have increasingly focused on improving the temporal reasoning capabilities of LLMs. Temporal reasoning in natural language processing (NLP) typically falls into three categories: temporal expression detection and normalization Li et al. (2024b), temporal relation extraction, and event forecasting Yuan et al. (2024). The integration of temporal graphs has enabled LLMs to perform more effectively in tasks such as time comparison Xiong et al. (2024). Several temporal QA datasets have been developed to test LLMs’ temporal reasoning abilities, including TEMPLAMA Dhingra et al. (2022); Tan et al. (2023), TemporalWiki Jang et al. (2022), and time-sensitive QA datasets Chen et al. (2021). By combining LLMs with temporal KGs Lee et al. (2023); Yuan et al. (2024); Xia et al. (2024), more accurate event forecasting has become possible.
@@
3 Beyond Single Pass, Looping Through Time: KG-IRAG Framework
The KG-IRAG framework, marked as $F$, is defined in equation 1, as a combination of 2-Agent Temporal Anchoring Module ($\mathcal{A}$) and the Iterative Reasoning Controller ($\mathcal{C}$). Let $q$ be the input query requiring temporal reasoning, the framework operates through sequential state transitions guided by knowledge graph traversals. Where $D^{(t)}=\Psi(S^{(t)})$ represents the triplets retrieved at iteration $t$ through the retrieval operator $\Psi$, with $D^{(0)}$ being the initial anchor triplets determined by LLM1. The main notations of KG-IRAG are summarized in Table 1, Appendix A.
$\mathcal{F}=(\mathcal{A},\mathcal{C}),\quad\mathcal{F}(q,D)=\mathcal{C}\left(\mathcal{A}(q,D^{(0)}),\{D^{(t)}\}_{t=0}^{T}\right)$
3.1 Formalization of Iterative Process
At the initial stage $t^{(0)}$, LLM1 generates initial retrieval parameters $(\tau^{(0)},\ell^{(0)})$ as well as reasoning prompts to be used in $\mathcal{C}$ through whole iteration, initial $\mathcal{A}$ is defined as
$\mathcal{A}(q,D^{(0)})=(\tau^{(0)},\ell^{(0)},\mathcal{P}^{(0)})=\underset{(\tau,\ell)\in\mathbb{T}\times\mathcal{L}}{\arg\max}\ P_{\text{LLM1}}(q|D_{\tau,\ell})$
The reasoning state at step(time) $t$ is defined as:
$S^{(t)}=(\tau^{(t)},\ell^{(t)},\mathcal{R}^{(t)},\mathcal{P}^{(t)})$ (3)
Where Temporal anchor point $\tau^{(t)}\in\mathbb{R}^{+}$; Spatial location context $\ell^{(t)}\in\mathcal{L}$; $\mathcal{R}^{(t)}\subset D$ stands for Retrieved triplets $\{(e_{s},r,e_{o})\}$ and $\mathcal{P}^{(t)}$ for Reasoning prompt template. The state evolution follows the Markovian transition:
$S^{(t+1)}=\Phi\left(S^{(t)},\Delta^{(t)}\right)$ (4)
$\Delta^{(t)}=\left(\mathcal{M}_{1}(q,S^{(t)}),\mathcal{M}_{2}(q,S^{(t)})\right)$ (5)
where $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ represent LLM1 and LLM2 respectively, with LLM1 $\mathcal{M}_{1}(q,S^{(t)})$ evaluates termination condition $\mathbb{I}_{\text{stop}}\in\{0,1\}$ based on LLM2 Generation, and LLM2 $\mathcal{M}_{2}(q,S^{(t)})$ generates new retrieval parameters $(\tau^{(t+1)},\ell^{(t+1)})$, controller $\mathcal{C}$ is defined as:
$\mathcal{C}(\{D^{(t)}\})=\mathbb{I}_{\text{stop}}\cdot\mathcal{M}_{\text{final}}\left(\bigoplus_{t=0}^{T}\mathcal{R}^{(t)}\right)+(1-\mathbb{I}_{\text{stop}})\cdot\mathcal{F}(q,D^{(t+1)})$
Algorithm 1 Iterative Reasoning Process
1:Query $q$, KG $D$, LLMs $\mathcal{M}_{1},\mathcal{M}_{2}$
2:Answer $ans$
3:$S^{(0)}\leftarrow\mathcal{A}(q,D^{(0)})$triangleright Initial anchoring
4:for $t=0$ to $T_{\text{max}}$ do
5:     $\mathcal{R}^{(t)}\leftarrow\Psi(S^{(t)},D)$triangleright Subgraph retrieval
6:     $\mathbb{I}_{\text{stop}}\leftarrow\mathcal{M}_{2}(\mathcal{R}^{(t)},\mathcal{P}^{(t)})$
7:     if $\mathbb{I}_{\text{stop}}=1$ then
8:         break
9:     end if
10:     $S^{(t+1)}\leftarrow\Phi(S^{(t)},\Delta^{(t)})$triangleright State transition
11:end for
12:$ans\leftarrow\mathcal{M}_{1_{\text{final}}}(\bigoplus_{t=0}^{T}\mathcal{R}^{(t)})$
13:return $ans$
3.2 Knowledge Graph Construction
The temporal knowledge graph can be formalized as a directed multigraph $D=(\mathcal{V},\mathcal{E})$ where: Vertex set $\mathcal{V}=\mathcal{E}_{\text{time}}\cup\mathcal{E}_{\text{loc}}\cup\mathcal{E}_{\text{event}}$ comprising time points, locations, and events; Labeled edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{R}\times\mathcal{V}$ with relation types =$\mathcal{R}=\{r_{\text{temp}},r_{\text{spatial}},r_{\text{causal}}\}$. The graph construction process follows the operator $D=\bigcup_{i=1}^{n}\Gamma(d_{i})$ where raw data points $d_{i}$ are transformed through:
$\Gamma(d_{i})=\{(t_{i},r_{\text{temp}},l_{i}),(l_{i},r_{\text{spatial}},e_{i}),(t_{i},r_{\text{causal}},e_{i})\}$
3.3 Triplet Retrieval
Once the starting time $\tau^{(0)}$ and location $\ell^{(0)}$ have been identified by LLM1, KG exploration begins. The retrieval process at step $t$ is governed by the operator $\Psi^{(t)}:D\times S^{(t)}\rightarrow\mathcal{R}^{(t)}$:
$\mathcal{R}^{(t)}=\bigcup_{k=1}^{K^{(t)}}\left\{(e_{s},r,e_{o})\in N_{k}(v^{(t)})\right\}$ (8)
Where Current spatiotemporal anchor $v^{(t)}=(\tau^{(t)},\ell^{(t)})$, and $N_{k}(v)$ is k-hop neighborhood expansion. The neighborhood expansion follows dynamic hop calculation:
$K^{(t)}=\left\lceil 2-\log_{2}\left(1+\frac{|\mathcal{R}^{(t-1)}|}{|\mathcal{R}^{(0)}|}\right)\right\rceil$ (9)
3.4 Answer Generation
Algorithm 2 Multi-Iteration Answer Synthesis
1:$\{\mathcal{R}^{(t)}\}_{t=0}^{T}$, $\mathbb{I}_{\text{stop}}$, $\mathcal{M}_{1_{\text{final}}}$, $\kappa$
2:Final answer $ans$
3:$\mathbf{E}\leftarrow\emptyset$
4:if $\mathbb{I}_{\text{stop}}=1$ then
5:     for $t=0$ to $T$ do
6:         $\mathbf{E}\leftarrow\mathbf{E}\cup\{\mathcal{R}^{(t)}\}$
7:     end for
8:     if $|\mathbf{E}|<\kappa$ then
9:         $ans\leftarrow\mathcal{M}_{\text{final}}(\mathbf{E})$triangleright Evidence Fusion
10:     else
@@ 3.4 Answer Generation
12:     end if
13:else
[…]
15:end if
16:return $ans$
Once the triplets from the various iterations is sufficient based on $\mathbb{I}_{\text{stop}}$, they are combined into a coherent RAG Data Prompt, and the sufficient triplet evidences are then processed together with the former reasoning prompt, together to get the answer generated. If data is insufficient to yield the answer or is out of boundary during KG lookup, ‘no answer’ is returned as the result. Let $\kappa$ be the threshold for maximum retrieval, the final answer synthesizes multi-step evidence through Eq 10, while algorithm 2 shows the detailed process for answer synthesis.
$ans=\begin{cases}\mathcal{M}_{1_{\text{final}}}\left(\sum_{t=0}^{T}\alpha^{(t)}\mathcal{R}^{(t)}\right)&\text{if }\sum_{t=0}^{T}|\mathcal{R}^{(t)}|<\kappa\\ \text{"no answer"}&\text{otherwise}\end{cases}$
3.5 Why NOT Single-Pass Retrieval?
Single-pass RAG must enlarge its temporal neighbourhood to at least the dependency depth $\hat{d}$ (Def. B.1) in a $\delta$-regular graph, yielding a cost of $\Theta(\delta^{\hat{d}})$. As formally shown in Lemma B.2 (Appendix B.1), the probability of retrieving the full evidence path will decays exponentially to $P_{\text{success}}=\delta^{-(\hat{d}-k)}$. Considering the complex in reasoning, single-pass retrieval faces an accuracy–efficiency dilemma: either pay an exponential cost or accept an exponentially small success rate. On the other hand, KG-IRAG issues $T=\lceil\hat{d}/k_{0}\rceil$ that focuses queries of radius $k_{0}\!\ll\!\hat{d}$, keeping cost $\Theta(\delta^{k_{0}})$ per-step while preserve higher accuracy (Thm. B.3).
3.6 Case Study
To further explain the details of ‘iteration‘ in KG-IRAG, a case study was conducted. Presented in section C, it simulate an initial travel plan to Sydney Opera House, including defining whether the initial plan is valid, as well as how to define an optimal time to adjust the trip.
4 Dataset properties
4.1 Data properties
To further test a RAG model’s ability to explore and do logical reasoning on temporal data, 3 datasets, weatherQA-Irish, weatherQA-Sydney and trafficQA-TFNSW are prepared for experiments. The construction of these datasets is tailored to allow the model to perform both entity-based retrieval and time-dependent reasoning. While the two weather datasets are used to identifying abnormal events like rainfall, the TFNSW dataset includes extensive numerical data on traffic volume, adding complexity to the questions by requiring numerical comparisons. All three datasets are transferred into KGs that capture location relationships and temporal records. While the details of datasets are presented in Appendix D, Table 2 presents attributes used to form the KGs, with time treated as an entity to enhance retrieval capabilities. Table 3 represents the size of each dataset and the corresponding KGs generated.
4.2 Question-Answer Generation
The design of the questions in the QA datasets is structured to progressively increase the complexity of the retrieval tasks. The first question (Q1) is a fundamental entity recognition and retrieval task, while the second (Q2) and third (Q3) questions introduce logical reasoning by incorporating time-dependent queries. The goal is to examine the system’s ability to not only retrieve relevant data but also engage in iterative reasoning over time. Table 4 shows the details of question proportion in the three datasets.
For each KG $D$, after anchor times $\mathcal{T}_{\text{anchor}}$ are sampled uniformly across $D$’s temporal coverage, question-answer pairs $\mathcal{Q}=\{(q_{k},a_{k})\}$ are generated through controlled synthesis:
$\mathcal{Q}=\bigcup_{t\in\mathcal{T}_{\text{anchor}}}\left(\mathcal{Q}_{\text{base}}(t)\cup\mathcal{Q}_{\text{reason}}(t)\right)$ (11)
4.2.1 Base Question(Q1): Abnormal Event Detection
Q1 is designed as a straightforward retrieval task in which the model identifies whether an "abnormal" event, such as rainfall or traffic congestion, occurred during a specific time slot. For each $t\sim\mathcal{T}_{\text{anchor}}$, generate Q1 as followed and answer derived in Equation 12:
“Considering I am (traveling/ on trip…) at [location] during [t, t+$\Delta$ t], consider [event restriction], can I avoid [abnormal event]? ”
$a_{\text{base}}(t)=\mathbb{I}\left(\exists e\in D\cap[t,t+\Delta t]\right)$ (12)
4.2.2 Questions 2 and 3: Leave Early or Late
In contrast to Q1, the design of Q2 and Q3 introduces temporal reasoning components that significantly increases the complexity of the retrieval process. Q2 asks the model to calculate the latest time at which one can leave early to avoid an abnormal event, while Q3 focuses on determining the earliest time to leave late. Both questions require the model to infer optimal departure times by reasoning over a temporal range, making use of the temporal relationships embedded in KGs. For positive Q1 cases, extend to temporal reasoning:
$q_{\text{early}}(t)$: “Considering same event length $\Delta t$, what’s the latest departure before $t$ to avoid [abnormal event]? ”
$q_{\text{late}}(t)$: “Considering same event length $\Delta t$, what’s the earliest departure after $t$ to avoid [abnormal event]? ”
With $L$ set as maximum retrieval duration (12 h / 9 h), the answers are computed using temporal constraint solving. Algorithm 3 present the overall procedure of QA generation.
$a_{\text{early}}(t)=\max\{t^{\prime}<t\ |\ \nexists e\in[t^{\prime},t^{\prime}+\Delta t],t^{\prime}>t-L\}$
$a_{\text{late}}(t)=\min\{t^{\prime}>t\ |\ \nexists e\in[t^{\prime},t^{\prime}+\Delta t],t^{\prime}<t+L\}$
Algorithm 3 QA Pair Generation
1:KG $D$, Max attempts $M$, Trip duration $\Delta t$, Max Retrieval Length $L$
2:Question-answer pairs $\mathcal{Q}$
[…]
4:for each $t\in\mathcal{T}_{\text{anchor}}$ do
5:     Generate $q_{1},a_{1}$ via Eq 12
6:     $\mathcal{Q}\leftarrow\mathcal{Q}\cup\{(q_{1},a_{1})\}$
7:     if $a_{1}=\text{True}$ then